OpenStudy (anonymous):

A transverse wave on a string is described by the following wave function. y = 0.080 sin(π/10x + 5πt) where x and y are in meters and t is in seconds. a) Determine the transverse speed at t = 0.210 s for an element of the string located at x = 1.80 m. b) Determine the transverse acceleration at t = 0.210 s for an element of the string located at x = 1.80 m. c) What is the wavelength of this wave? d) What is the period of this wave? e) What is the speed of propagation of this wave?

OpenStudy (ybarrap):

The easy stuff first: \( \omega=5\pi\): The angular frequency \(\large T={2\pi\over\omega}\) : The period \(k=\dfrac \pi {10}\): The wave number \(\large \lambda=\large {2\pi\over k}\): The wavelength \(\nu=\large { \lambda \over T}\) or \(\nu=\large { \omega \over k}\): The speed of Propagation To determine the transverse velocity and acceleration, we need to find \(\dfrac {dy}{dx}\) and \(\dfrac {d^2y}{dx^2}\) and then substitute the values for t and x that are given. Hope this helped. Good luck.